# Physical models of embryonic epithelial healing: A review

### v1

@@ Introduction
Embryonic wound healing in epithelial monolayers refers to the rapid repair of tissue gaps during early development, essential to preserve its integrity and barrier function. Unlike adult wound healing, it is different in several ways: it is faster, does not leave scars, has minimized inflammatory response and has more frequent junctional rearrangements. In contrast, adult wound healing involves the recruitment of immune cells, matrix deposition by fibroblasts and scarring. With these differences, embryonic wound repair is a minimal system to study tissue repair driven primarily by epithelial dynamics. It is, therefore, ideal to test mechanistic models that aim to predict tissue closure from cellular and subcellular dynamics. Wound closure is achieved through two primary, often coexisting, mechanisms: a purse-string around the wound edge and crawling of epithelial cells into the gap. While crawling is common to both embryonic and adult healing, the purse-string is more prevalent in embryonic healing and generally less prevalent in adult epithelia [1, 2, 3, 4, 5].
@@
Both mechanisms depend on time-dependent cytoskeletal remodelling, adhesive dynamics and coupling between force generation and cell polarity. The purse-string requires spatially varying models of junctional tension and contractility, while crawling requires models of dynamic cell polarity and interactions with the extracellular matrix. Both are regulated by mechanical and biochemical cues, including epidermal growth factor (EGF), calcium influx (Ca2+), reactive oxygen species (ROS) and extracellular adenosine tri-phosphate (ATP) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 2, 43, 44, 45].
@@
  • Cell-based models - The observed tissue behavior emerges from interactions among multiple cells, at different levels of detail. These include lattice models, network models, particle models and phase field models;
@@
Neither approach is inherently superior, but they serve different purposes depending on the research question. Cell-based models are useful for simulating collective cell migration observed in embryonic healing from individual cell interactions. Continuum models can represent large-scale deformations and mechanical properties of epithelial tissues during wound closure providing insight on the stresses and strains the tissue experiences during this process.
@@
  • Mechanical assumptions - vital for accurately modeling contractile forces, as cells must deform and adhere to one another while responding to mechanical stresses:
@@
    • Epithelial tissues form confluent monolayers that are continuous sheets without gaps;
@@
Implementing all these characteristics into a single model is not feasible, so simplifications and abstractions are fairly common. As the focus of our work will be on epithelial monolayers, these results may not be immediately generalizable to other tissues [53]. Connective tissues lack the tight cell junctions which are characteristic of epithelial tissues, with their interactions primarily mediated by the ECM [54]. Muscle fibers have contractile properties and spatial organization that differs significantly from epithelial sheets [55]. Neural tissues are made of neurons which have a unique morphology with axons and dendrites, with interactions mediated by synapses, rather than continuous junctions [56].
@@
In addition to the previously mentioned categories of assumptions, we may include additional properties in embryonic models. Polarity is an example of such a property, as a relevant factor in many embryonic healing processes [43]. There are several polarity mechanisms that can be explored, but they can be grouped into cell-autonomous mechanisms [57, 58, 59] and non-cell-autonomous mechanisms [60, 61, 62, 63]. Choosing one mechanism constrains both the biological regimes applicable and the model’s validity domain. There may be distinctions between leader cells and non-leader cells, particularly in embryonic healing processes, and different types of cells may have different responses to some signals [49].
@@
Collective motion in cells may be driven by several factors, one of them being chemical gradients, in a process described as chemotaxis. Chemical signaling during chemotaxis is usually modeled by linking extracellular chemical gradients to intracellular signaling pathways that steer cell motion. Cells can produce their own chemoattractants and degrade them to create local gradients, resulting in cell aggregation and self-organized clusters. Meanwhile, the self-generated chemical gradients enable long-range coordination of cell movements by linking local chemoattractant production to global migration dynamics [64, 65, 66]. Hybrid models, commonly applied in cancer invasion studies, also provide insight into embryonic healing, bridging tissue-scale interactions, and cellular microenvironments.
One of the challenges in modeling tissue mechanics is extending theories developed in 2D to three-dimensional contexts. This is especially relevant when scaling models from individual cells to full tissue behavior and from in vitro settings to in vivo conditions [51, 67]. Our review mainly address results for models in 2D, however there are experimental observations which show 3D models may be more accurate to address apico-basal polarity’s role during healing [253].
@@
Cellular packing also plays a central role in regulation of tissue growth, material properties and cell communication. Discrete models are very useful to understand these dynamics [67]. Cell-based models are particularly useful when describing relatively small tissues, where finite-size effects and fluctuations become relevant to the observed dynamics, being particularly useful in microscopic phenomena [50].
@@
The Cellular Potts Model (CPM), also known as Glazier-Graner-Hodwege model, represents cells or subcellular components as clusters of lattice sites that can change their configuration based on energy minimization principles [69]. Cell rearrangement in CPM is probabilistic and modeled using the Boltzmann distribution, which reflects energy changes resulting from interactions between neighboring sites while respecting constraints on volume or area, controlled through a temperature parameter. The system’s energy $E$ is typically given by:
@@
CPM has been widely used to model cell sorting, tissue growth and cell movement within tissues, as well as morphogenesis [74, 75, 76, 77, 78, 79, 80, 81]. While lattice models are computationally efficient and allow the observation of emergent tissue behavior, it is not without drawbacks. The discretization of space may lead to artifacts in simulations, by ignoring off-lattice behaviors. Additionally, the temperature parameter lacks a clear experimental analog. And finally, it does not fully address cell-substrate interactions [73, 51, 82].
@@
The most widely used network model is the vertex model, where the tissues are typically confluent, and the vertices set the degrees of freedom of the system. The positions of these vertices evolve over time according to forces arising from mechanical interactions between neighboring cells. The energy function is typically of the form:
@@
Continuum models are an alternative way to describe collective cell behavior on scales where individual cell details are less relevant than average quantities like density, velocity, or chemical concentrations and their evolution under biological and physical constraints. These models treat tissues as continuous media instead of relying on discrete structural units and capture emergent behaviors of cell ensembles [50].
@@
We start by looking at phase field models as they are at the interface between cell-based and continuum tissue models. Phase field models are widely used to simulate dynamic interfaces in tissues because they represent boundaries implicitly, removing the need to track their positions explicitly. In these models, a phase field or order parameter $\phi$ transitions smoothly from 1 inside the tissue to 0 outside, allowing for a continuous representation of tissue structure [49].
@@
Here, the first term accounts for interfacial energy, while the second term incorporates cell-cell adhesion, surface tension, and tissue elasticity, often modelled by a Landau potential energy $f(\phi)=a(1-\phi^{2})^{2}$, where $a$ is a proportionality constant, ensuring $f$ has units of energy. The parameter $\epsilon$ controls the interface width, smoothing sharp transitions [128, 129]. In the Allen-Cahn formulation (based on the Ginzburg-Landau equation), or dissipative dynamics, the phase field evolves according to:
@@
A more detailed version of this framework is the multiphase field approach, which models tissues as collections of interacting phases, each representing an individual cell. Instead of modeling smooth boundary dynamics, they explicitly describe interactions between phases, using conservation laws and enforcing total area and/or volume conservation through additional constraints. A common formulation for the free energy in this case is:
@@
In hydrodynamic tissue models, cells are no longer individualized. Instead, they describe tissues as viscous or viscoelastic fluids, giving a framework to study tissue flow, stress distribution and large-scale deformations using principles from classical hydrodynamics and associated computational tools. Unlike passive fluids, tissues exhibit active stresses driven by cell motility and contractile forces, which must be incorporated into the hydrodynamic equations [48, 50].
@@
The left side of the equation corresponds to the material or convective derivative, while the right side includes the influence of the pressure field $p(\vec{\mathbf{r}},t)$ and of viscosity (coupled to a dynamic viscosity coefficient $\eta$). $\sigma^{\text{A}}$ is a stress tensor typically dependent on the local orientation or polarity of cells, reflecting internal contractile forces such as actomyosin activity (e.g. $\sigma^{\text{A}}=\zeta\mathbf{Q}$, where $\zeta$ is an activity coefficient that controls the magnitude of active stress, and $\mathbf{Q}$ is the nematic tensor, which we will discuss further below). In biological tissues, inertial effects are typically negligible due to low Reynolds numbers, allowing for simplification to overdamped or Stokes-flow regimes, where $\vec{\mathbf{v}}\cdot\vec{\nabla}\vec{\mathbf{v}}$ can be ignored, in many practical applications [146]. This makes hydrodynamic models particularly useful in wound closure studies, where large-scale collective flows dominate. Below we discuss two particular classes of hydrodynamic models that are relevant to modelling tissue dynamics: active polar and nematic models.
@@
Hydrodynamic, polar and nematic models have been extensively applied to understand the mechanisms of embryonic healing and cell migration. Early hydrodynamic models demonstrated how monolayer elasticity, adhesive tension, and lamellipodial crawling interact during closure, revealing that partial closure can occur even without lamellipodia [33]. Further studies highlighted the role of gap shape and area in determining closure speed, with crawling aiding closure while adhesion hinders it [155]. More recent hydrodynamic frameworks have studied the influence of extracellular signal-regulated kinase (ERK) concentration fields on cell migration [156]. A continuum model linking single-cell behavior to collective migration showed that the activity of PIEZO1, a mechanosensitive ion channel, inhibits leader cell formation and suppresses directionality, emphasizing its inhibitory role in reepithelialization [157].
@@
In continuum mechanics, we distinguish between material or Lagrangian coordinates $\mathbf{X}=X_{I}\mathbf{E}_{I}$, which label each material point in the initial undeformed configuration $B_{0}$, and spatial or Eulerian coordinates $\mathbf{x}=x_{i}e_{i}$, which label the same point at the current deformed configuration $B_{t}$. The two coordinates are related to each other through a mapping $\mathbf{x}=f(\mathbf{X},t)$, which can be linearized to describe stretching and rotations through the deformation tensor $\mathbf{F}=\frac{\partial\mathbf{x}}{\partial\mathbf{X}}$. In this way, we are able to define constitutive laws in reference to either the original or deformed configuration, as seen in the following sections.
@@
Viscoelastic models capture both short-term elastic recovery and long-term viscous flow, closely mirroring how biological tissues spread and recover shape under sustained forces.

### v2

@@ Introduction
Embryonic wound healing in epithelial monolayers refers to the rapid repair of tissue gaps during early development, essential to preserve its integrity and barrier function. Unlike adult wound healing, it is different in several ways: it is faster, does not leave scars, has minimized inflammatory response and has more frequent junctional rearrangements. In contrast, adult wound healing involves the recruitment of immune cells, matrix deposition by fibroblasts and scarring. With these differences, embryonic wound repair is a minimal system to study tissue repair driven primarily by epithelial dynamics. It is, therefore, ideal to test mechanistic models that aim to predict tissue closure from cellular and subcellular dynamics. Wound closure is achieved through two primary, often coexisting, mechanisms: a purse-string around the wound edge and crawling of epithelial cells into the gap (see Figure 1). While crawling is common to both embryonic and adult healing, the purse-string is more prevalent in embryonic healing and generally less prevalent in adult epithelia [1, 2, 3, 4, 5].
@@
Both mechanisms depend on time-dependent cytoskeletal remodelling, adhesive dynamics and coupling between force generation and cell polarity. The purse-string requires spatially varying models of junctional tension and contractility, while crawling requires models of dynamic cell polarity and interactions with the extracellular matrix. Both are regulated by mechanical and biochemical cues, including epidermal growth factor (EGF), calcium influx (Ca2+), reactive oxygen species (ROS) and extracellular adenosine tri-phosphate (ATP). Laser ablation experiments have also been used to probe embryonic tissue mechanics in vivo and constrain models of epithelial wound closure [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 2, 43, 44, 45, 46].
@@
In addition to the purse-string and crawling mechanisms, closure of certain excisional wounds in multilayered embryonic epithelia, such as the Xenopus ectoderm, is also driven by contraction and ingression of cells exposed within the wound, forming characteristic ‘Y’-shapes [47, 48].
@@
In line with this focus, we prioritize studies where these models are applied explicitly to embryonic epithelial wound healing. For some models, wound-specific applications remain limited. In those cases, we also discuss studies using the same physical models to address closely related epithelial processes, provided that the modeled mechanisms are directly relevant to wound closure or that the original works explicitly note their applicability to healing geometries. Such inclusions are used selectively and only when necessary to represent a given modeling approach. Throughout, we restrict our discussion to epithelial monolayers, while drawing connections to morphogenetic studies only where they directly inform healing mechanisms.
The embryonic epithelial systems covered in this review include laser-induced wounds in the Drosophila pupal notum, epidermis and imaginal disc, with occasional consideration of other embryonic systems such as chicks and zebrafish (Danio rerio). For studies that do not directly address embryonic healing, the corresponding systems involve in vitro epithelial monolayers subjected to scratch assays or gap-closure geometries. These epithelial monolayers studied include, most commonly, MDCK (Madin-Darby canine kidney) and HaCaT (immortalized human keratinocyte) cells. These systems serve as physical models of epithelial mechanics and collective behavior relevant to wound healing, rather than as models of embryological development in the strict sense.
@@
  • Cell-based models - The observed tissue behavior emerges from interactions among multiple cells, at different levels of detail. These include lattice models, network models and phase field models;
@@
Neither approach is inherently superior, but they serve different purposes depending on the research question. Cell-based models are useful for simulating collective cell migration observed in embryonic healing from individual cell interactions. Continuum models can represent large-scale deformations and mechanical properties of epithelial tissues during wound closure providing insight on the stresses and strains the tissue experiences during this process (see Table 1).
@@
  • Cell mechanical assumptions - vital for accurately modeling contractile forces, as cells must deform and adhere to one another while responding to mechanical stresses:
@@
    • Epithelial tissues form confluent monolayers - continuous sheets without holes;
@@
Implementing all these characteristics into a single model is not feasible, so simplifications and abstractions are fairly common. As the focus of our work will be on epithelial monolayers, these results may not be immediately generalizable to other tissues [56]. Connective tissues lack the tight cell junctions which are characteristic of epithelial tissues, with their interactions being primarily mediated by the ECM, while epithelial cells have persistent junctional mechanics that couple the cells, and the ECM is interacted with at the basal face [57]. Muscle fibers have contractile properties and spatial organization that differs significantly from epithelial sheets [58]. Neural tissues are made of neurons which have a unique morphology with axons and dendrites, with interactions mediated by synapses, rather than continuous junctions [59].
@@
Considering the assumptions outlined above, we focus on multicellular embryonic wounds, which are treated physically as free boundaries within a contiguous, confluent epithelial monolayer. In addition to the previously mentioned categories of assumptions, we may include additional properties in embryonic models. Polarity is an example of such a property, as a relevant factor in many embryonic healing processes [43]. There are several polarity mechanisms that can be explored, but they can be grouped into cell-autonomous mechanisms [60, 61, 62] and non-cell-autonomous mechanisms [63, 64, 65, 66]. Choosing one mechanism constrains both the biological regimes applicable and the model’s validity domain. There may be distinctions between leader cells and non-leader cells, particularly in embryonic healing processes, and different types of cells may have different responses to some signals [52].
@@
Collective motion in cells may be driven by several factors, one of them being chemical gradients, in a process described as chemotaxis. Chemical signaling during chemotaxis is usually modeled by linking extracellular chemical gradients to intracellular signaling pathways that steer cell motion. Cells can produce their own chemoattractants and degrade them to create local gradients, resulting in cell aggregation and self-organized clusters. Meanwhile, the self-generated chemical gradients enable long-range coordination of cell movements by linking local chemoattractant production to global migration dynamics [67, 68, 69]. Hybrid models, commonly applied in cancer invasion studies, also provide insight into embryonic healing, bridging tissue-scale interactions, and cellular microenvironments. One of the challenges in modeling tissue mechanics is extending theories developed in 2D to three-dimensional contexts. This is especially relevant when scaling models from individual cells to full tissue behavior and from in vitro settings to in vivo conditions [54, 70]. Our review mainly address results for models in 2D, however there are experimental observations which show 3D models may be more accurate to address apico-basal polarity’s role during healing. A clear example of this limitation is for cell-ECM junctions, which occur at the basal surface, and can result in out-of-plane deformations. Most of the considered models incorporate ECM through effective interactions. Additionally, effects of cells layers below are not considered, which fails to capture non-canonical closure mechanisms [47, 256].
Model Schematic Features Scale Strengths Limitations
Lattice [Uncaptioned image] Discrete lattice representation; Cells as collection of pixels; Stochastic updates via energy minimization; Cell motion via pixel rearrangements. Subcellular to cellular Efficient for large cell populations; Handles cell rearrangements naturally; Can include heterogeneity and stochasticity. Interactions propagate discretely; Timescale is algorithm-dependent; Discretizations limit mechanical accuracy at tissue scale.
Network [Uncaptioned image] Cells represented as polygons; Displacement of vertices (Vertex) via energy minimization; Displacement of centers (Voronoi) via energy minimization with geometry reconstructed; Junctional forces and contractility encoded along edges; Confluence constraint . Cellular Explicitly captures junctional and anisotropic tension; Handles cell rearrangements; Efficient for confluent epithelia. Coarse-grained subcellular dynamics; Cannot resolve smooth boundary deformations (protrusions); Nonconfluent tissues are not naturally represented in the models.
Phase-field [Uncaptioned image][Uncaptioned image] Multiphase Cells are continuous scalar fields; Diffuse interface between cells; Cell shape and motion via field dynamics. Binary Tissue and environment separation via two scalar fields; No explicit individual cells; Interface motion via Cahn-Hilliard/Allen-Cahn dynamics. Cellular (Multiphase) and Tissue (Binary) Multiphase Smooth cell boundaries; Cell deformation, growth and division emerge from continuous geometric dynamics; Can simulate many cells without explicit computational tracking. Binary Efficient for large-scale tissue dynamics; Captures wound interface motion and large scale instabilities; Can include active stresses. Multiphase High computational cost for many cells; Interface thickness must be tuned carefully; Junctional mechanics are phenomenological rather than explicit. Binary Cannot resolve single-cell mechanics; Junction-level and subcellular processes coarse-grained; Not suitable to study anisotropy or heterogeneity at cellular scale.
Hydro-dynamic [Uncaptioned image] Tissue-scale continuum scalar and vector fields; Flows in epithelial sheets via force-balance equations; Symmetry choice may be isotropic, polar or nematic. Tissue Handles large-scale collective flows and stress propagation Naturally includes active stresses Can include polarity alignment Can describe both isotropic and anisotropic tissue behaviors Subcellular details are not resolved Assumes continuum approximation Loses cellular heterogeneity
Visco-elastic [Uncaptioned image] Tissue treated as a viscoelastic material; Constitutive stress-strain relation; Time-dependent stress relaxation and material response; Tissue Simple framework to capture bulk tissue mechanics Directly relates stress and strain Easily coupled to external forces or boundary conditions Does not capture active stresses Subcellular processes are absent Does not naturally include anisotropic or orientational order
Table 1: Table summarizing the main models explored in the review. The listed features correspond to the minimal structural elements of each model class; additional couplings and fields commonly used in the literature are discussed in the main text.
@@
These models are coarse-grained at the cellular level, where internal cell dynamics are abstracted away into effective behaviors. These models assume that tissue mechanics is dominated by cell-level processes such as adhesion, cortical tension, and neighbour rearrangements, with tissue-scale behavior emerging from their collective organization. Cellular packing also plays a central role in regulation of tissue growth, material properties and cell communication. Discrete models are very useful to understand these dynamics [70]. Cell-based models are particularly useful when describing relatively small tissues, where finite-size effects and fluctuations become relevant to the observed dynamics, being particularly useful in microscopic phenomena [53].
@@
Fig 2: Schematic of a lattice model of wound healing, where $\mathcal{W}$ is the wound region and the different colored pixel regions correspond to different cells.
The Cellular Potts Model (CPM), also known as Glazier-Graner- Hogeweg model, represents cells or subcellular components as clusters of lattice sites that can change their configuration based on energy minimization principles [72]. Cell rearrangement in CPM is probabilistic and modeled using the Boltzmann distribution, which reflects energy changes resulting from interactions between neighboring sites while respecting constraints on volume or area, controlled through a temperature parameter (see Figure 2). The system’s energy $E$ is typically given by:
@@
CPM has been widely used to model cell sorting, tissue growth and cell movement within tissues, as well as morphogenesis [77, 78, 79, 80, 81, 82, 83, 84]. While lattice models are computationally efficient and allow the observation of emergent tissue behavior, it is not without drawbacks. The discretization of space may lead to artifacts in simulations, by ignoring off-lattice behaviors and strain propagation occurs discretely, limiting mechanical accuracy at tissue-scale. Additionally, the temperature parameter lacks a clear experimental analog, and the implicit notion of time in the Metropolis algorithm makes dynamic interpretation difficult. While extensions such as kinetic Monte Carlo can approximate timescales, they require knowledge of event rates that are typically unknown. And finally, it does not fully address cell-substrate interactions [76, 54, 85].
@@
Fig 3: Schematic of a network model of wound healing, where $\mathcal{W}$ is the wound region and the different colored polygons correspond to cells with different number of neighbours.
The most widely used network model is the vertex model, where the tissues are typically confluent, and the vertices set the degrees of freedom of the system (see Figure 3). The positions of these vertices evolve over time according to forces arising from mechanical interactions between neighboring cells. The energy function is typically of the form:
@@
Continuum models are an alternative way to describe collective cell behavior on scales where individual cell details are less relevant than average quantities like density, velocity, or chemical concentrations and their evolution under biological and physical constraints. Tissues are treated as a material with effective mechanical properties like elasticity, viscosity or active stress, rather than a collection of cells[53].
@@
We start by looking at phase field models as they are at the interface between cell-based and continuum tissue models. Phase field models are widely used to simulate dynamic interfaces in tissues because they represent boundaries implicitly, removing the need to track their positions explicitly (see Figure 4). In these models, a phase field or order parameter $\phi$ transitions smoothly from 1 inside the tissue to 0 outside, allowing for a continuous representation of tissue structure [52].
Fig 4: (Left) Schematic of a phase field model of wound healing where $\mathcal{W}$ is the wound region where $\phi=0$ and the tissue region is where $\phi=1$. (Right) Schematic of a multiphase field model of wound healing where $\phi_{i}$ is the phase field associated with cell ii.
@@
Here, the first term accounts for interfacial energy, while the second term incorporates cell-cell adhesion, surface tension, and tissue elasticity, often modelled by a Landau potential energy $f(\phi)=a(1-\phi^{2})^{2}$, where $a$ is a proportionality constant, ensuring $f$ has units of energy density. The parameter $\epsilon$ controls the interface width, smoothing sharp transitions [131, 132]. In the Allen-Cahn formulation (based on the Ginzburg-Landau equation), or dissipative dynamics, the phase field evolves according to:
@@
A more detailed version of this framework is the multiphase field approach, which models tissues as collections of interacting phases, each representing an individual cell. This is effectively a cell-based model. Instead of modeling smooth boundary dynamics, they explicitly describe interactions between phases, using conservation laws and enforcing total area and/or volume conservation through additional constraints. A common formulation for the free energy in this case is:
@@
In hydrodynamic tissue models, cells are no longer individualized. Instead, they describe tissues as viscous or viscoelastic fluids, giving a framework to study tissue flow, stress distribution and large-scale deformations using principles from classical hydrodynamics and associated computational tools (see Figure 5). Unlike passive fluids, tissues exhibit active stresses driven by cell motility and contractile forces, which must be incorporated into the hydrodynamic equations [51, 53].
@@
The left side of the equation corresponds to the material or convective derivative, while the right side includes the influence of the pressure field $p(\vec{\mathbf{r}},t)$ and of viscosity (coupled to a dynamic viscosity coefficient $\eta$). $\sigma^{\text{A}}$ is a stress tensor typically dependent on the local orientation or polarity of cells, reflecting internal contractile forces such as actomyosin activity (e.g. $\sigma^{\text{A}}=\zeta\mathbf{Q}$, where $\zeta$ is an activity coefficient that controls the magnitude of active stress, and $\mathbf{Q}$ is the nematic tensor, which we will discuss further below). In biological tissues, inertial effects are typically negligible due to low Reynolds numbers, allowing for simplification to overdamped or Stokes-flow regimes, where $\vec{\mathbf{v}}\cdot\vec{\nabla}\vec{\mathbf{v}}$ can be ignored, in many practical applications [149].
Fig 5: Schematic of an active hydrodynamic model of wound healing. $\mathcal{W}$ corresponds to the wound region, and $\vec{\mathbf{u}}$ is the velocity field, driven by active stress (blue and red regions).
This makes hydrodynamic models particularly useful in wound closure studies, where large-scale collective flows dominate. Below we discuss two particular classes of hydrodynamic models that are relevant to modelling tissue dynamics: active polar and nematic models.
@@
Hydrodynamic, polar and nematic models have been extensively applied to understand the mechanisms of cell migration although not much use has been directed towards embryonic healing. Early hydrodynamic models demonstrated how monolayer elasticity, adhesive tension, and lamellipodial crawling interact during closure, revealing that partial closure can occur even without lamellipodia [33]. Further studies highlighted the role of gap shape and area in determining closure speed, with crawling aiding closure while adhesion hinders it [158]. More recent hydrodynamic frameworks have studied the influence of extracellular signal-regulated kinase (ERK) concentration fields on cell migration [159]. A continuum model linking single-cell behavior to collective migration showed that the activity of PIEZO1, a mechanosensitive ion channel, inhibits leader cell formation and suppresses directionality, emphasizing its inhibitory role in reepithelialization [160].
@@
In continuum mechanics, we distinguish between material or Lagrangian coordinates $\mathbf{X}=X_{I}\mathbf{E}_{I}$, which label each material point in the initial undeformed configuration $B_{0}$, and spatial or Eulerian coordinates $\mathbf{x}=x_{i}e_{i}$, which label the same point at the current deformed configuration $B_{t}$. The two coordinates are related to each other through a mapping $\mathbf{x}=f(\mathbf{X},t)$, which can be linearized to describe stretching and rotations through the deformation tensor $\mathbf{F}=\frac{\partial\mathbf{x}}{\partial\mathbf{X}}$ (see Figure 6). In this way, we are able to define constitutive laws in reference to either the original or deformed configuration, as seen in the following sections.
@@
Viscoelastic models capture both short-term elastic recovery and long-term viscous flow, closely mirroring how biological tissues spread and recover shape under sustained forces (see Figure 7).
Fig 7: Schematic of a viscoelastic model of wound healing. Where $\mathcal{W}$ is the wound region and displacement is driven by the internal tissue stress $\sigma$.